# Admissibility of Stein Shrinkage for Batch Normalization in the Presence of Adversarial Attacks

### v1

@@ 3 Dominance of Stein Shrinkage in the Presence of Adverserial Attacks
Consider the problem of estimating a parameter vector $\theta\in\mathbb{R}^{p}$ based on corrupted observations $Z=X+Y,$ where $X\sim\mathcal{N}_{p}(\theta,\sigma^{2}I_{p})$ and $Y\in\mathbb{R}^{p}$ is independent, mean-zero sub-Gaussian vector with a known proxy variance $2\varepsilon^{2}.$ Our goal is to compare the risks (expectation of the loss) of the naive estimator $\hat{\theta}^{\mathrm{MLE}}$ and the JS-estimator
@@
under squared error loss. It is well-known that $\hat{\theta}^{\mathrm{JS}}$ strictly dominates $\hat{\theta}^{\mathrm{MLE}}$ when $Y=0$ and $p\geq 3$ James-Stein61 . We aim to show that this dominance remains valid even under additive sub-Gaussian noise, provided the proxy variance $2\varepsilon^{2}$ is sufficiently small.
@@
Theorem 3.1.
Let $\theta\in\mathbb{R}^{p}$ be an unknown parameter vector. Suppose we observe $Z=X+Y$, where $X\sim\mathcal{N}_{p}(\theta,I_{p})$ and $Y\in\mathbb{R}^{p}$ is an independent, mean-zero sub-Gaussian vector with parameter $2\varepsilon^{2}$, i.e., for all unit vectors $v\in\mathbb{R}^{p}$ and all $\lambda\in\mathbb{R}$, $\mathbb{E}\left[\exp\left(\lambda\langle v,Y\rangle\right)\right]\leq\exp\left(\varepsilon^{2}\lambda^{2}\right)$. Define the JS-estimator by $\hat{\theta}^{\mathrm{JS}}=\left(1-\frac{p-2}{\|Z\|^{2}}\right)Z,$ and the maximum likelihood estimator by $\hat{\theta}^{\mathrm{MLE}}=Z$. Then, for all $\theta\in\mathbb{R}^{p}$ and for all $p\geq 3$ the JS-estimator strictly dominates the MLE in mean squared risk: $R(\hat{\theta}^{\mathrm{JS}},\theta)<R(\hat{\theta}^{\mathrm{MLE}},\theta).$
We now provide a sketch of the proof and refer the reader to the Appendix for more details.
Sketch of Proof.
We can rewrite the risk of the JS-estimator as:
$R(\hat{\theta}^{\mathrm{JS}},\theta)=R(\hat{\theta}^{\mathrm{MLE}})-2(p-2)\mathbb{E}\left(\frac{Z^{\top}(Z-\theta)}{Z^{\top}Z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{Z^{\top}Z}\right).$ (4)
@@
Further, by applying Cauchy–Schwarz and Jensen’s inequalities we prove (in the Appendix) that: $\mathbb{E}\left(\frac{2Z^{\top}\theta+p-2}{Z^{\top}Z}\right)<2.$ Hence, the risk of the James–Stein estimator is strictly less than that of the MLE for $p\geq 3$ completing the proof. ∎
@@ 3.2 Dominance of the JS-estimator under Sub-Gaussian Perturbations for Gamma Distributed Random Variables
In many applications, empirical variances follow a scaled chi-squared distribution, which is a special case of the gamma distribution. Given observations $x_{ij}$, define the empirical variance: $\hat{\sigma}_{i}^{2}=\frac{1}{n}\sum_{j=1}^{n}(x_{ij}-\bar{x}_{i})^{2}\sim\mathrm{Gamma}\left(\alpha=\frac{n-1}{2},\ \beta_{i}=\frac{2\sigma_{i}^{2}}{n}\right)$. Thus, estimating $\sigma_{i}^{2}$ reduces to estimating the scale parameter $\beta_{i}$ in a gamma distribution with fixed shape parameter $\alpha$. For this task, we can use the following theorem from literature:
Theorem 3.2 (JS-Estimator Shrinkage Dominance for the Gamma Model DasGupta86AOS ).
@@
Now, applying the above theorem to empirical variances, we obtain: $\hat{\sigma}_{i}^{2,\mathrm{JS}}=\frac{n}{n+1}\hat{\sigma}_{i}^{2}+cV,\quad V=\left(\prod_{j=1}^{p}\hat{\sigma}_{j}^{2}\right)^{1/p},\quad c\in\left[0,\frac{4n(p-1)}{(n+1)((n-1)p+2)}\right].$ We now present our second key result namely, that this dominance persists under additive sub-Gaussian noise for the Gamma distributed random variable.
Theorem 3.3.
[…]
Remark 3.4.
@@
[…]
Thus, the difference in risks depends on the sign of the second term in the above expression expressed as: $c\left(2\sum_{i=1}^{p}\mathbb{E}\left(\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right)+p\mathbb{E}\left(cV^{2}\right)\right).$ To show that this quantity is negative, we analyze the sign depending on whether $c>0$ or $c<0$, and whether the linear term dominates the quadratic term. We consider two main cases:
@@
    $02\sum_{i=1}^{D}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]>$ 0 (5)
@@
    $02\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]<$ 0 (6)
@@
Lemma 3.5 (Stein’s Lemma for the Gamma Distribution).
Let $X\sim\mathrm{Gamma}(\alpha,\beta)$, and let $h(x)$ be a differentiable function such that $\mathbb{E}[Xh(X)]<\infty\quad\text{and}\quad\mathbb{E}[Xh^{\prime}(X)]<\infty$. Then the following (Stein’s) identity holds: $\mathbb{E}[(X-\alpha\beta)h(X)]=\beta\,\mathbb{E}[Xh^{\prime}(X)]$.
We apply Stein’s identity to the function $h(x)=V$, where $V$ is the geometric mean defined by $V=\left[\prod_{i=1}^{p}\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}\right)\right]^{1/p},\quad X_{i}=\hat{\sigma}_{x,i}^{2}$. Using the Stein’s identity for the Gamma distributed random variable, which is given by, $\mathbb{E}[h(X)]=\frac{1}{\alpha\beta}\mathbb{E}[Xh(X)]-\frac{1}{\alpha}\mathbb{E}[Xh^{\prime}(X)],$ we differentiate $h(X)$ with respect to $\hat{\sigma}_{x,k}^{2}$ and get: $\frac{dV}{d\hat{\sigma}_{x,k}^{2}}=\frac{1}{p}\cdot V\cdot\left(\frac{1}{\hat{\sigma}_{x,k}^{2}+\hat{\sigma}_{y,k}^{2}+w_{k}}\right).$ Substituting into Stein’s identity for $X_{i}=\hat{\sigma}_{x,i}^{2}$, we get: $\mathbb{E}[V]=\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]-\frac{1}{\alpha p}\mathbb{E}\left[V\cdot\frac{\hat{\sigma}_{x,i}^{2}}{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}}\right].$
Depending on the sign and structure of the risk difference (as in the two main cases above), we proceed to estimate the relevant expectations using various inequalities — such as Bernstein’s inequality (to control sub-exponential terms like $W_{i}$), Jensen’s inequality, Cauchy–Schwarz, and the arithmetic–geometric mean inequality. These tools allow us to bound the expressions for a given choice of $c$ and derive the dominance for the JS-estimator of the variance parameter of the Normally distributed random variable under sub-Gaussian perturbations.
Further technical details and precise bounds are deferred to the appendix. ∎
@@ 4 Experiments
In this section, we present several experimental results demonstrating the performance of BN with our proposed shrinkage estimators in both image classification and segmentation tasks, under standard and adversarial settings. To simulate adversarial attacks, we added noise samples drawn from a sub-Gaussian distribution with a fixed variance. The sub-Gaussian density is given by $f(x)=\frac{1}{2\sqrt{2}\pi\sigma}\left(\frac{x^{2}}{2\sigma^{2}}+\frac{1}{4}\right)^{-1},$ derived from a mixture of Lévy and Gaussian distributions, and the samples were drawn using the inverse transform sampling technique. GEORGIOU20063061 . For classification, we employed the ResNet-9 he2015deepresiduallearningimage architecture on the CIFAR-10  cifar10 dataset, as well as a 3D CNN model on the PPMI dataset ppmi-data . For object segmentation, we used the HRNetV2 wang2020deep model applied to the Cityscapes datasetCordts2015Cvprw . Since BN is known to become unstable when the batch size is small, we report results for varying batch sizes and sub-Gaussian noise levels. Our goal is to demonstrate that our shrinkage-based method is more robust to adversarial perturbations and offers greater overall stability. We remind the reader that our goal here is not to improve the SOTA results on these data in the literature, but to demonstrate the performance gains from the use of BN in the presence (absence) of adversarial attacks with the correct use of Stein shrinkage compared to vanilla BN and other forms of shrinkage applied to BN. Throughout all the experiments, we avoided training-time tricks and focused on obtaining a honest comparison.
BN Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
@@
Our (32) 80.89 $\pm$ 0.51 60.7 $\pm$ 0.96 44.55 $\pm$ 1.4 35.34 $\pm$ 1.6 26.67 $\pm$ 3.38 25.66 $\pm$ 3.63
Our (64) 80.78 $\pm$ 0.54 61.7 $\pm$ 0.81 45.17 $\pm$ 1.43 36.19 $\pm$ 0.57 32.05 $\pm$ 0.58 25.26 $\pm$ 0.69
Our (128) 79.1 $\pm$ 0.3 60.49 $\pm$ 1.06 44.96 $\pm$ 1.3 36.27 $\pm$ 1.01 29.2 $\pm$ 4.67 28.12 $\pm$ 3.32
Khoshsirat khoshsirat2023improvingnormalizationjamessteinestimator (32) 76.82 $\pm$ 1.01 57. 49 $\pm$ 1.71 41.93 $\pm$ 1.46 33.20 $\pm$ 1.73 27.10 $\pm$ 3.5 25.39 $\pm$ 3.72
Khoshsirat khoshsirat2023improvingnormalizationjamessteinestimator (64) 79.84 $\pm$ 1.18 58.71 $\pm$ 1.48 43.75 $\pm$ 1.67 33.92 $\pm$ 2.4 30.75 $\pm$ 1.32 23.36 $\pm$ 1.20
Khoshsirat khoshsirat2023improvingnormalizationjamessteinestimator (128) 79.12 $\pm$ 0.27 60.02 $\pm$ 1.11 44.85 $\pm$ 1.47 36.35 $\pm$ 1.47 29.15 $\pm$ 4.25 28.34 $\pm$ 3.26
@@
Table 1: CIFAR-10 Test data classification performance using BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean accuracy $\pm$ standard deviation.
In the experiments, we compare our approach mainly with that of Khoshsirat et al. khoshsirat2023improvingnormalizationjamessteinestimator , aiming to show that theoretically correct estimation of mean as well as variance shrinkage significantly improves performance, particularly in small batch regimes. In addition, with the correct variance shrinkage, one can achieve higher accuracy in the presence of adversarial perturbations. We also tested a variant using only the mean shrinkage estimator and no shrinkage on variance. Additionally, we included models regularized with Lasso and Ridge methods for comparison, as these are commonly used shrinkage methods. Detailed formulas for these estimators and complete experimental setup are provided in the Appendix B. Note that in all the tables depicting the results, we have highlighted the best performing BN method (ours) for small batch sizes that are preferred from a computational efficiency perspective as well as when data are scarce.
4.1 Application to Cifar10 Data
We conducted experiments on the CIFAR-10 dataset, which consists of 50,000 training and 10,000 test images across 10 classes. We simulate sub-gaussian noise by perturbing the input features with zero-mean noise of varying magnitudes $(0-30\%)$, applied independently to each image during validation and testing. The results in Table 1 confirm that our proposed BN significantly outperforms the baselines in noisy settings across all batch sizes. Although standard BN variants quickly degrade under noise, our method maintains substantially higher accuracy, especially with moderate to large noise levels $(10-30\%).$ Compared to the method in khoshsirat2023improvingnormalizationjamessteinestimator , our approach shows consistent improvement due to the correct development and use of variance shrinkage. The method with only mean correction (see Appendix) performs worse than our full model, highlighting the importance of variance correction. Finally, Lasso and Ridge-based approaches do not improve noise performance, showing behavior similar to the standard BN.

### v2

@@ 3 Dominance of Stein Shrinkage in the Presence of Adversarial Attacks
Consider the problem of estimating a parameter vector $\theta\in\mathbb{R}^{p}$ based on corrupted observations $Z=X+Y,$ where $X\sim\mathcal{N}_{p}(\theta,\sigma^{2}I_{p})$ and $Y\sim\mathrm{SG}(2\varepsilon^{2})$. Our goal is to compare the risks (expectation of the loss) of the naive estimator $\hat{\theta}^{0}=Z$ and the JS-estimator
@@
under squared error loss. It is well known that $\hat{\theta}^{\mathrm{JS}}$ strictly dominates $\hat{\theta}^{0}$ when $Y=0$ and $p\geq 3$ (James and Stein, 1961). We aim to show that this dominance remains valid even under additive sub-Gaussian noise, provided the proxy variance $2\varepsilon^{2}$ is sufficiently small.
@@
Theorem 1
Let $\theta\in\mathbb{R}^{p}$ be an unknown parameter vector. Suppose that we observe $Z=X+Y$, where $X\sim\mathcal{N}_{p}(\theta,I_{p})$ and $Y\in\mathbb{R}^{p}$ is an independent $Y\sim\mathrm{SG}(2\varepsilon^{2})$. Define the JS-estimator by $\hat{\theta}^{\mathrm{JS}}=\left(1-\frac{p-2}{\|Z\|^{2}}\right)Z,$ and the observation-based estimator by $\hat{\theta}^{0}=Z$. Then, for all $\theta\in\mathbb{R}^{p}$ and for all $p\geq 3$ the JS-estimator strictly dominates the observation-based estimator in mean squared risk: $R(\hat{\theta}^{\mathrm{JS}},\theta)<R(\hat{\theta}^{\mathrm{0}},\theta).$
We now provide a sketch of the proof for this theorem and refer the reader to the Appendix A for a detailed proof.
Proof [Sketch of the Proof] We can rewrite the risk of the JS-estimator as:
$R(\hat{\theta}^{\mathrm{JS}},\theta)=R(\hat{\theta}^{\mathrm{0}})-2(p-2)\mathbb{E}\left(\frac{Z^{\top}(Z-\theta)}{Z^{\top}Z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{Z^{\top}Z}\right).$ (4)
@@
Further, by applying Cauchy–Schwarz and Jensen’s inequalities we prove (in the Appendix) that:
$\mathbb{E}\left(\frac{2Z^{\top}\theta+p-2}{Z^{\top}Z}\right)<2.$
Hence, the risk of the JS-estimator is strictly less than that of the observation-based estimator for $p\geq 3$, thus completing the proof.
@@ 3.2 Dominance of the JS-estimator under Sub-Gaussian Perturbations for Gamma Distributed Random Variables
In many applications, SVs follow a scaled chi-squared distribution, which is a special case of the gamma distribution. Given observations $x_{ij}$, define the SV: $\hat{\sigma}_{i}^{2}=\frac{1}{n}\sum_{j=1}^{n}(x_{ij}-\bar{x}_{i})^{2}\sim\mathrm{Gamma}\left(\alpha=\frac{n-1}{2},\ \beta_{i}=\frac{2\sigma_{i}^{2}}{n}\right)$. Thus, estimating $\sigma_{i}^{2}$ reduces to estimating the scale parameter $\beta_{i}$ in a Gamma distribution with fixed shape parameter $\alpha$. For this task, we can use the following theorem from literature (Anirban, 1986):
Theorem 2 (Dominance of the JS-Estimator for the Gamma Model)
Let $X_{1},\dots,X_{p}\sim\mathrm{Gamma}(\alpha,\beta_{i})$ be independently distributed. Consider the naive estimator $\hat{\beta}_{i}^{\,0}=\frac{X_{i}}{\alpha+1},i=1,\dots,p.$ Define the JS-type shrinkage estimator $\hat{\beta}_{i}^{\,\mathrm{JS}}=\hat{\beta}_{i}^{\,0}+cV,V=\Biggl(\prod_{j=1}^{p}X_{j}\Biggr)^{\!1/p}.$ If $c\in\left[\,0,\;\frac{2(p-1)}{(\alpha+1)(\alpha p+1)}\,\right],$ then the shrinkage estimator dominates the naive estimator in risk, i.e. $R\!\left(\hat{\beta}^{\,\mathrm{JS}},\beta\right)\;\leq\;R\!\left(\hat{\beta}^{\,0},\beta\right),\forall\;\beta\in\mathbb{R}_{+}^{p}.$
Applying the previous theorem to SVs, we obtain the JS-type shrinkage estimator
$\hat{\sigma}_{i}^{2,\,\mathrm{JS}}\;=\;\frac{n}{n+1}\,\hat{\sigma}_{i}^{2}\;+\;c\,V,\qquad V\;=\;\Biggl(\prod_{j=1}^{p}\hat{\sigma}_{j}^{2}\Biggr)^{\!1/p},\qquad c\;\in\;\left[\,0,\;\frac{4n(p-1)}{(n+1)\big((n-1)p+2\big)}\,\right].$
We now present our second key result, namely, that this dominance persists even under additive sub-Gaussian noise for the Gamma-distributed random variables.
Theorem 3
@@
Remark 4
@@
$R(\hat{\beta}^{\,\mathrm{JS}},\beta)=\mathbb{E}\left(\left\|\hat{\beta}^{\,\mathrm{JS}}-\beta\right\|_{2}^{2}\right)=\mathbb{E}\left(\left\|\hat{\beta}^{\,0}+cV-\beta\right\|_{2}^{2}\right)=$
$=\mathbb{E}\left(\left\|\hat{\beta}^{\,0}-\beta\right\|_{2}^{2}\right)+2\sum_{i=1}^{p}\mathbb{E}\left(\left(\hat{\beta}_{0,i}-\beta_{i}\right)\cdot cV\right)+p\mathbb{E}\left(c^{2}V^{2}\right)$
Thus, the difference in risks depends on the sign of the second term in the above expression expressed as:
$c\left(2\sum_{i=1}^{p}\mathbb{E}\left(\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right)+p\mathbb{E}\left(cV^{2}\right)\right).$
To show that this quantity is negative, we analyze the sign depending on whether $c>0$ or $c<0$, and whether the linear term dominates the quadratic term. We consider two main cases:
@@
    $2\sum_{i=1}^{D}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]>0$ (5)
@@
    $2\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]<0$ (6)
@@
Lemma 5 (Stein’s Lemma for the Gamma Distribution)
Then the following (Stein’s) identity holds:
$\mathbb{E}[(X-\alpha\beta)h(X)]=\beta\,\mathbb{E}[Xh^{\prime}(X)].$
We apply Stein’s identity to the function $h(x)=V$, where $V$ is the geometric mean defined by $V=\left[\prod_{i=1}^{p}\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}\right)\right]^{1/p},\quad X_{i}=\hat{\sigma}_{x,i}^{2}$. Using the Stein’s identity for the Gamma distributed random variable, which is given by,
we differentiate $h(X)$ with respect to $\hat{\sigma}_{x,k}^{2}$ and get: $\frac{dV}{d\hat{\sigma}_{x,k}^{2}}=\frac{1}{p}\cdot V\cdot\left(\frac{1}{\hat{\sigma}_{x,k}^{2}+\hat{\sigma}_{y,k}^{2}+w_{k}}\right).$ Substituting into Stein’s identity for $X_{i}=\hat{\sigma}_{x,i}^{2}$, we get:
Depending on the sign and structure of the risk difference (as in the two main cases above), we proceed to estimate the relevant expectations using various inequalities — such as Bernstein’s inequality (to control sub-exponential terms like $W_{i}$), Jensen’s inequality, Cauchy–Schwarz, and the arithmetic–geometric mean inequality respectively. These tools allow us to bound the expressions for a given choice of $c$ and derive the dominance for the JS-estimator of the variance parameter of the normally distributed random variable under sub-Gaussian perturbations. Further technical details and precise bounds are deferred to the appendix.
@@ 4 Experiments
In this section, we present several experimental results demonstrating the performance of BN with our proposed shrinkage estimators in both image classification and segmentation tasks, under standard and adversarial settings. To simulate adversarial attacks, we generated samples of the noise signal modeled by a sub-Gaussian–like distribution with the density
$f(x)=\frac{1}{2\sqrt{2}\pi\sigma}\left(\frac{x^{2}}{2\sigma^{2}}+\frac{1}{4}\right)^{-1},$
which arises from a mixture of Lévy and Gaussian distributions (Panayiotis G. and Chris, 2006). The samples were drawn using inverse transform sampling.
For classification, we employed the ResNet-9 (He et al., 2016) architecture on the CIFAR-10 dataset (Alex et al., 2009), as well as a 3D CNN model on the PPMI dataset (1). For object segmentation, we used the HRNetV2 (Wang et al., 2020) model applied to the Cityscapes dataset (Cordts et al., 2015). Since BN is known to become unstable when the batch size is small, we report results for varying batch sizes and sub-Gaussian noise levels. Our goal is to demonstrate that our shrinkage-based method is more robust to adversarial perturbations and offers greater overall stability.
We remind the reader that our goal here is not to improve the SOTA results on these data in the literature, but to demonstrate the performance gains from the use of BN in the presence (absence) of adversarial attacks with the correct use of Stein shrinkage compared to vanilla BN and other forms of shrinkage applied to BN. Throughout all experiments, we avoided training-time tricks and focused on obtaining an honest comparison.
In the experiments, we compare our approach mainly with Khoshsirat and Kambhamettu (2024), aiming to show that theoretically correct estimation of mean as well as variance shrinkage significantly improves performance, particularly in small batch regimes. In addition, with the correct variance shrinkage, one can achieve higher accuracy in the presence of adversarial perturbations. We also tested a variant using only the mean shrinkage estimator and no shrinkage on variance. Additionally, we included models regularized with Lasso and Ridge methods for comparison, as these are classical shrinkage methods in literature. Below we summarize the forms of these shrinkage estimators and provide explicit formulas for convenience. Derivations can be found, for example, in Hastie et al. (2009).
Lasso mean and variance shrinkage. These estimators are obtained by applying an $\ell_{1}$ penalty in penalized least squares problems (Hastie et al. (2009)). For the mean, this leads to the soft-thresholding solution $\hat{\mu}=\operatorname{sign}(\bar{x})\max\!\left(0,|\bar{x}|-\tfrac{\lambda}{2n}\right),$ where $\bar{x}=\tfrac{1}{n}\sum_{i=1}^{n}x_{i}$ is the SM. This form encourages the shrinkage of small mean values towards zero, potentially setting them exactly to zero when $\lambda$ is large. For variance, an analogous penalized formulation yields $\hat{\sigma}^{2}=\max\!\left(0,\,s^{2}-\tfrac{\lambda}{2}\right),$ where $\lambda>0$ is the regularization hyperparameter and $s^{2}=\tfrac{1}{n}\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}$ is the SV.
Ridge mean and variance shrinkage. These estimators are obtained by applying an $\ell_{2}$ penalty in penalized least squares problems (Hoerl and Kennard (1970); Hastie et al. (2009)). For the mean, this yields $\hat{\mu}=\frac{\sum_{i=1}^{n}x_{i}}{n+\lambda},$ and for the variance $\hat{\sigma}^{2}=\frac{s^{2}}{1+\lambda},$ where $s^{2}=\tfrac{1}{n}\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}$ is the SV and $\lambda>0$ is the regularization parameter.
Complete experimental setup are provided in the Appendix B. Note that in all the tables depicting the results, we have highlighted the best performing BN method (ours) for small batch sizes that are preferred from a computational efficiency perspective as well as when data are scarce. Code available at https://github.com/sivolgina/shrinkage.
4.1 Application to Cifar10 Data
We conducted experiments on the CIFAR-10 dataset, which consists of 50,000 training and 10,000 test images across 10 classes. We simulate sub-gaussian noise by perturbing the input features with zero-mean noise of varying magnitudes $(0-30\%)$, applied independently to each image during validation and testing. The results in Table 1 confirm that our proposed BN significantly outperforms the baselines in noisy settings across all batch sizes. Although standard BN variants quickly degrade under noise, our method maintains substantially higher accuracy, especially with moderate to large noise levels $(10-30\%).$ Compared to the method in Khoshsirat and Kambhamettu (2024), our approach shows consistent improvement due to the correct development and use of variance shrinkage. The method with only mean correction performs worse than our full model, highlighting the importance of variance correction. Finally, Lasso and Ridge-based approaches do not improve noise performance, showing behavior similar to the standard BN.
Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
@@ 4.1 Application to Cifar10 Data
Ours (32) 80.89 $\pm$ 0.51 60.7 $\pm$ 0.96 44.55 $\pm$ 1.4 35.34 $\pm$ 1.6 26.67 $\pm$ 3.38 25.66 $\pm$ 3.63
Ours (64) 80.78 $\pm$ 0.54 61.7 $\pm$ 0.81 45.17 $\pm$ 1.43 36.19 $\pm$ 0.57 32.05 $\pm$ 0.58 25.26 $\pm$ 0.69
Ours (128) 79.1 $\pm$ 0.3 60.49 $\pm$ 1.06 44.96 $\pm$ 1.3 36.27 $\pm$ 1.01 29.2 $\pm$ 4.67 28.12 $\pm$ 3.32
Mean corrected (32) 78.7 $\pm$ 1.32 56.79 $\pm$ 0.87 40.13 $\pm$ 1.19 31.57 $\pm$ 1.19 25.82 $\pm$ 3.65 24.77 $\pm$ 3.97
Mean corrected (64) 80.6 $\pm$ 0.4 54.15 $\pm$ 1.26 38.24 $\pm$ 0.66 30.17 $\pm$ 1.37 26.78 $\pm$ 1.42 20.85 $\pm$ 1.13
Mean corrected (128) 79.9 $\pm$ 0.48 49.2 $\pm$ 1.35 32.96 $\pm$ 0.99 25.97 $\pm$ 1.2 25.82 $\pm$ 3.65 24.77 $\pm$ 3.97
Khoshsirat et al. (32) 76.82 $\pm$ 1.01 57. 49 $\pm$ 1.71 41.93 $\pm$ 1.46 33.20 $\pm$ 1.73 27.10 $\pm$ 3.5 25.39 $\pm$ 3.72
Khoshsirat et al. (64) 79.84 $\pm$ 1.18 58.71 $\pm$ 1.48 43.75 $\pm$ 1.67 33.92 $\pm$ 2.4 30.75 $\pm$ 1.32 23.36 $\pm$ 1.20
Khoshsirat et al. (128) 79.12 $\pm$ 0.27 60.02 $\pm$ 1.11 44.85 $\pm$ 1.47 36.35 $\pm$ 1.47 29.15 $\pm$ 4.25 28.34 $\pm$ 3.26
@@
Table 1: Performance of ResNet-9 with different BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean accuracy $\pm$ standard deviation on test CIFAR-10 data. Numbers inside parentheses in the far left column denote batch size.